Please help with the problem from a work sheet.
Four times the sum of two consecutive integers equals 2 more than 14 times the larger integer. Find the integers.

Answer by ptaylor(1894)

Let x=the larger integer
Then x-1=the smaller integer
Four times the sum=4(x+x-1)=4(2x-1)
14 times the larger integer=14(x)
2 more than 14 times the larger integer=14x+2
Now we are told that
4(2x-1)=14x+2 get rid of parens
8x-4=14x+2 subtract 14x from and also add 4 to both sides
8x-14x-4+4=14x-14x+2+4 collect like terms
-6x=6 divide both sides by -6
x=-1-----------------------------------larger integer
x-1=-1-1=-2-----------------------------smaller integer
CK
4(-3)=14(-1)+2
-12=-14+2
-12=-12

Hope this helps-----ptaylor